bianca07:

A force of 73.8 newtons is applied to travel bag at an angle 39.0° with the horizontal. What are the horizontal and vertical components of the force

Narad:

Let the force be x N at an angle of \[\theta \] with the horizontal The horizontal component is \[=x *\cos \theta \]and the vertical component is \[x *sin \theta \]